I just need help with part 2. Use the graphical method to construct the shear-force and bending-moment diagrams for the beam shown. Let a=4.0 ft, b=9.0 ft,*Part 2 Your answer is partially correct. Try again. Determine the shear force acting at each of the following locations: (a)
Answer #1

e a eft bgft c5ft B с 10 d = eft W = 8Kips P = 75 kips FBD P FRE Ra Ra+Ring = 75-68119) 2. taking moment about a (RB)(4+9+5+4a) Vata=0 = - 30.54 Kips. by vat 2-457 : -30.54 kips. e) v atx: 1357 (-30.54 +849) 41.46 Kips - d) rat 2=18$t (Just to left o